a man spends Rs 5 on bar chocolates. had they been 2 and a half paisa, he would have bought 10 more. how many did he buy?
Answer by richwmiller(17219)

Assuming they were Rs5 for each bar.
n*5=2.5(n+10)
He originally bought 10 bars at Rs5 each
check
5*10=2.5*(10+10)
50=2.5*20
ok